OpenStudy (xbeenchantedx):

Solve the equation below. Check for extraneous solutions. sqrt(x+7) = x+5

OpenStudy (xbeenchantedx):

\[\sqrt{x+7} = x + 5\]

OpenStudy (anonymous):

square both sides first

OpenStudy (anonymous):

\[(\sqrt{x+7})^{2} = (x+ 5)^{2}\]

OpenStudy (anonymous):

distribute (x+5) ^2 \[x+ 7 = x^{2} + 10x + 25\] \[x^{2} +9x +18 = 0\] Now factor this expression. Find x.

OpenStudy (xbeenchantedx):

Thank you!